The air in a bottle is at atmospheric pressure 1.01 x 105 Pa and a temperature of 300C. At temperature 990C the cork is ejected out from the bottle. What was the pressure in the bottle just before the cork left?
Answer #1

T1 = 300+273.15 = 573.15K

T2 = 990+273.15 = 1263.15K

P1 = 1.01×105pa

At constant volume ,. P1/T1 = P2/T2

=) P2 = 1.01 ×105 × 1263.15/573.15 = 2.226×105 pa